# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ 2.3 Survey Operations
In §2.3.1 we describe main-survey operations, which DESI focused on almost exclusively in the time period covered by DR1, 2021 May 14 through 2022 June 13. Observations of SV tiles were included in the EDR and are described in DESI Collaboration et al. (2024b, a), although we summarize these observations briefly in §2.3.2 since they are being re-released in DR1. Finally, §2.3.3 briefly introduces the set of observations taken as part of the special survey included in DR1.
2.3.1 The Main Survey
The DESI main survey consists of 9929 dark tiles, 5676 bright tiles, and 2657 backup tiles, which DESI aims to observe over the course of its 5-year survey (Schlafly et al., 2023).888As the DESI survey is performing better than anticipated, the dark- and bright-time programs have been extended—both in terms of the number of tiles and the exact footprint—since the start of science operations. However, in this paper we adopt the survey definition in place at the start of the survey, as documented in Schlafly et al. (2023), since it is that survey that was used in the DESI year-one cosmology analyses. For the dark- and bright-time surveys, the locations of the centers of these tiles are chosen to cover the Galactic caps accessible to DESI, roughly the sky where Legacy Surveys imaging is available (Dey et al., 2019), supplemented with modest additional cuts to remove tiles at low Galactic latitude (roughly ; see Schlafly et al. 2023). Meanwhile, the backup program footprint covers a wider declination range (roughly ) and lower Galactic latitude (roughly ), as backup-program targets are selected from Gaia photometry (Gaia Collaboration et al., 2016) which is available over the whole sky (Myers et al. 2023a; Dey et al. 2025, in preparation).
| No. of | No. of | No. of | Effective TimeaaThe effective exposure time is the on-sky integration time in reference, or ideal conditions (see §2.1 and Guy et al. 2023 for details). | AreabbThe area covered by tiles is larger than the true effective area available to targets due to bright star exclusions, focal plane geometry, hardware configuration, and higher priority targets blocking lower priority targets. | |
@@ 2.3.1 The Main Survey
Although the tile centers are fixed, the actual assignment of targets to fibers is not defined until we observe a particular location on the sky. At that time, the Merged Target List (MTL; Schlafly et al., 2023) is used to assign the highest-priority target to each fiber, in order to produce a tile with a fixed fiber-assignment configuration. The primary motivation for this “on the fly” design strategy is to be able to identify and increase the priority of quasars from DESI spectra so that they will be repeatedly observed. At , the Ly forest redshifts into the DESI spectral coverage (see §2.1), making these targets especially valuable probes of large-scale structure at high redshifts. However, because we do not know in advance which quasars are at , we are forced to learn this from the DESI spectra themselves. Once Ly-forest QSOs have been identified, we update the MTL with the results to ensure that future DESI observations will observe these targets whenever possible. The MTL allows us to identify times when observations of targets failed, for example, due to being assigned to a positioner which has stopped functioning. Observations of these targets may then be repeated in the future.
The need to use information about past observations in order to inform future observations leads to the following survey operations mode or sequence: {enumerate*}
@@
The DESI spectroscopic pipeline processes each night’s observations as they come in, and usually redshifts are available the following morning (see §3.1 and Guy et al., 2023).999The DESI team informally and fondly refers to this process as “redshifts by breakfast.” Following a quality assurance process by members of the survey operations team, the reduced data are incorporated into the MTL, roughly twice per week (Schlafly et al., 2023).
Tiles are selected for observation each night automatically by the Next Tile Selector (NTS; Schlafly et al., 2023). The basic scheme is to prefer equatorial fields while minimizing slewing and obtaining the minimum airmass possible. The NTS avoids observing areas of the sky where observations have been made that have not been incorporated into the MTL, so that any remaining Ly quasars in those regions can be identified. The NTS is also responsible for deciding whether to observe bright, dark, or backup tiles. This decision is made on the basis of the current survey speed, which is based on real-time measurements of the transparency, seeing, and sky background from the Exposure Time Calculator (ETC; Kirkby et al., 2025). The survey speed is expressed as a fraction relative to nominal dark conditions, where the sky background is 21.07 mag arcsec-2 in the -band, the seeing is 11, and conditions are photometric. When the survey speed is greater than 0.4 we observe the dark program; when the survey speed is greater than 0.08 but less than 0.4 we observe the bright program; otherwise we observe backup program tiles.
@@
2.3.2 Survey Validation
Although the primary focus of DR1 is on the new main-survey data, DR1 also includes all the SV data which were taken before the start of the main survey: {enumerate*}
SV1 (Target Selection Validation) tiles were used to verify and refine the target-selection algorithms for the main survey, and include dedicated secondary-target tiles which were used for special programs before we developed the special survey (see §2.3.3).
@@ 2.3.3 Special Tiles
3 Data Reduction and Data Products
In this section we describe the major DR1 data products and how they are produced. First, in §3.1 we present a brief overview of how the two-dimensional spectra are reduced to derive redshifts, redshift-quality flags, and spectral classifications. Next, in §3.2 we summarize the number of unique redshifts in DR1 and show the sample distributions in redshift and on the sky. In §3.3 we outline the contents and organization of the data products included in DR1, and in §3.4 we provide a brief introduction to the large-scale structure (LSS) catalogs used to carry out the DESI cosmological analysis and describe where and how those catalogs can be accessed. Finally, in §3.5 we describe the current set of value-added catalogs (VACs) accompanying DR1.
3.1 From Raw Spectra to Redshifts
The following three subsections describe the data-reduction pipeline (§3.1.1), the determination of redshifts and spectral classifications (§3.1.2), and the improvements in the spectroscopic pipeline relative to the data released in the EDR (§3.1.3).
3.1.1 Spectral Extraction and Calibration
@@
Using these data, we derive a master bias frame for each CCD, the spectral trace coordinates, wavelength calibration parameters, two-dimensional point-spread functions (PSFs), and a flat-field correction for each fiber of each camera. The processing of a scientific exposure acquired during the subsequent night proceeds as follows: We first pre-process each CCD image by subtracting the bias, the overscan level, and the dark current. We mask pixels affected by cosmic-ray hits and CCD defects, convert counts into electrons, and estimate the noise per pixel. We then adjust the coordinates of the spectral traces and the wavelength calibration of each fiber using the known wavelengths of the emission lines from the night-sky spectrum, while offsetting this solution such that the extracted spectra will be in a vacuum Solar System barycentric frame. These coordinate corrections are typically smaller than a CCD pixel relative to the afternoon calibrations. Next, we extract one-dimensional spectra using the spectroperfectionism algorithm of Bolton & Schlegel (2010). The results are counts per fiber and wavelength on a common wavelength grid (vacuum, barycentric), along with the resolution matrix, which quantifies the effective instrument resolution of each extracted spectrum as a function of wavelength. The spectra are then flat-fielded and sky-subtracted. Next, we use observations of F-type stars to determine the instrument throughput and to convert the measured counts into calibrated spectral energy density.111111All extracted (one-dimensional) DESI spectra are in units of . Finally, depending on which data products are being generated (see §3.3), multiple observations of the target or TILEID (see §2.3) are coadded using optimal (inverse variance) weights.
@@
We record the resolution matrix for each spectrum as a band-diagonal sparse matrix. A perfect resolution model should be multiplied by this matrix to achieve the equivalent observed model at the effective resolution of DESI. The underlying mathematics and motivation for the resolution matrix are described in §3 of Bolton & Schlegel (2010) and in §4.5 and Appendices C and D of Guy et al. (2023). We refer interested readers to the repository of DESI tutorials (see §4.3) for worked examples of how to use the resolution matrix in scientific analyses.
@@
For each spectrum, DESI also measures the template signal-to-noise ratio squared (TSNR2) for each of its major target classes BGS, LRG, ELG, QSO, and LYA.121212Here, LYA refers to QSOs which have been spectroscopically confirmed to be Ly forest () quasars. Also note that we do not derive a TSNR2 value for MWS targets. These quantities measure a wavelength-averaged, squared, signal-to-noise ratio, where the noise is determined from the current observations (CCD noise, sky level), the signal amplitude depends on the flux calibration, and the variation of the signal with wavelength is designed to best predict the redshift measurement precision of a given target class. For example, TSNR2_ELG gives larger weight to redder wavelengths covering the [O II] doublet at the redshifts of typical DESI ELG targets, while TSNR2_QSO gives more weight to bluer wavelengths important for measuring QSO emission lines. TSNR2_LRG is scaled to derive the effective exposure time (described in §2.3) of DESI dark-time exposures, while TSNR2_BGS is used to derive the effective exposure time for bright-time tiles, though this is also normalized to the reference dark conditions (see §4.14 of Guy et al. 2023 for more details).
3.1.2 Redshifts and Spectral Classifications
@@
With the fully calibrated spectra in-hand, we use Redrock to determine the optimum redshift and spectral classification of each object (Ross et al. 2020; Anand et al. 2024, Bailey et al. 2025, in preparation). Briefly, Redrock fits principal component analysis (PCA) templates of three broad, independent classes of objects—stars, galaxies, and quasars, corresponding to spectral type STAR, GALAXY, and QSO, respectively—on a spectral type-dependent grid of redshift. The fit which produces the lowest overall value yields the best-fitting redshift (Z), redshift uncertainty (ZERR), spectral type (SPECTYPE), and template coefficients (COEFF). In addition, Redrock reports the (minimum) of the best fit (CHI2) and (DELTACHI2), which is the difference between the two lowest values across all three spectral classes. Interpreted using Gaussian statistics, DELTACHI2 represents the statistical significance of the best fit relative to the next best fit; for example, a DELTACHI2 value of 25 implies that the best fit is “better” than the next best fit (not accounting for systematic uncertainties). Finally, besides redshifts and classification, a key metric reported by Redrock is the per-object redshift warning bitmask, ZWARN. In essence, a value of ZWARN=0 indicates that there are no known problems with either the input spectroscopic data nor with the corresponding redshift fit. We refer the reader to the DESI data model131313https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#zwarn, which documents the full set of possible ZWARN bits and their definitions, and to §3.2 of DESI Collaboration et al. (2024b) for more discussion.
@@
Finally, the DESI pipeline carries out three additional processing steps: emission-line fitting (emlinefit; Raichoor et al., 2023); a fit to the Mg II doublet (Chaussidon et al., 2023); and a quasar neural-network classifier called QuasarNET (Busca & Balland 2018; Farr et al. 2020; Green et al. 2025, in preparation), all three of which are described in more detail in §3.1.4 of DESI Collaboration et al. (2024b).141414Colloquially, DESI team members refer to these steps of the pipeline as “after-burners,” since they are collaborator-contributed algorithms executed after the original core pipeline Redrock results have been written out. Detailed analyses show that these after-burners are essential for evaluating successful ELG redshifts and for correctly classifying and measuring the correct redshifts for some classes of QSOs (Alexander et al., 2023; Lan et al., 2023; Raichoor et al., 2023; Chaussidon et al., 2023; DESI Collaboration et al., 2024a), so the after-burner catalogs are included as standard pipeline outputs.
| Parameter | Value | ReferenceaaReferences G23, DESI24c, and K24 are Guy et al. (2023), DESI Collaboration et al. (2024c), and Koposov et al. (2024), respectively. |
@@
3.1.3 Pipeline Improvements and Algorithmic Performance
Algorithmic updates for the Iron production in DR1 (relative to the Fuji specprod used in the EDR and Guadalupe; see §1.2) include new QSO templates split by redshift range (Brodzeller et al., 2023); improved sky subtraction by modeling the measured fiber throughput as a function of positioner location151515https://github.com/desihub/desispec/pull/1801 which replaces the simpler per-fiber, per-exposure measured normalization described in §4.7.2 of Guy et al. (2023); and multiple smaller bug-fixes. In addition, compared to the Fuji release in the EDR, most intermediate pipeline files are now gzipped to save disk space, although the coadded spectra and measured catalog files remain uncompressed for faster read-access (see §3.3).
@@ 3.1.3 Pipeline Improvements and Algorithmic Performance
3.2 Sample Distributions
After just 13 months of science operations, DESI has measured confident redshifts for approximately 18.7M unique science targets across all surveys and programs, making DR1 the largest sample of extragalactic redshifts ever assembled. In the main survey alone, DR1 includes a total of 8.5M, 9.0M, and 1.2M objects with reliable redshifts from the bright, dark, and backup programs, respectively. Here and in the rest of this section, we use the (logical “and”) condition to select a sample of unique objects with confident redshifts. The criterion identifies objects with no known hardware, observing, data quality, or redshift-fitting flags (see §3.1), and ZCAT_PRIMARY chooses the “best” observation of objects which have been observed in different surveys or programs (see §3.3). For additional details regarding the full range of possible data-quality bit-masks, see the DESI data model.161616https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#bit-masks-in-desi Alternatively, most users of the DESI data will likely be interested in the tracer-dependent quantitative criteria used to define the samples used for DESI cosmological analyses, which are described in §4.2 of DESI Collaboration et al. (2024b) and DESI Collaboration et al. (2024c) for the EDR and DR1 cosmological results, respectively.

### v2

@@ II.3 Survey Operations
In §II.3.1 we describe main-survey operations, which DESI focused on almost exclusively in the time period covered by DR1, 2021 May 14 through 2022 June 13. Observations of SV tiles were included in the EDR and are described in DESI Collaboration et al. (2024b, a), although we summarize these observations briefly in §II.3.2 since they are being re-released in DR1. Finally, §II.3.3 briefly introduces the set of observations taken as part of the special survey included in DR1.
II.3.1 The Main Survey
The DESI main survey consists of 9929 dark tiles, 5676 bright tiles, and 2657 backup tiles, which DESI aims to observe over the course of its 5-year survey (Schlafly et al., 2023).888As the DESI survey is performing better than anticipated, the dark- and bright-time programs have been extended—both in terms of the number of tiles and the exact footprint—since the start of science operations. However, in this paper we adopt the survey definition in place at the start of the survey, as documented in Schlafly et al. (2023), since it is that survey that was used in the DESI year-one cosmology analyses. For the dark- and bright-time surveys, the locations of the centers of these tiles are chosen to cover the Galactic caps accessible to DESI, roughly the sky where Legacy Surveys imaging is available (Dey et al., 2019), supplemented with modest additional cuts to remove tiles at low Galactic latitude (roughly ; see Schlafly et al. 2023). Meanwhile, the backup program footprint covers a wider declination range (roughly ) and lower Galactic latitude (roughly ), as backup-program targets are selected from Gaia photometry (gaia-collaboration16a) which is available over the whole sky (Myers et al., 2023; dey25a).
| No. of | No. of | No. of | Effective TimeaaThe effective exposure time is the on-sky integration time in reference, or ideal conditions (see §II.1 and Guy et al. 2023 for details). | AreabbThe area covered by tiles is larger than the true effective area available to targets due to bright star exclusions, focal plane geometry, hardware configuration, and higher priority targets blocking lower priority targets. | |
@@ II.3.1 The Main Survey
Although the tile centers are fixed, the actual assignment of targets to fibers is not defined until we observe a particular location on the sky. At that time, the Merged Target List (MTL; Schlafly et al., 2023) is used to assign the highest-priority target to each fiber, in order to produce a tile with a fixed fiber-assignment configuration. The primary motivation for this “on the fly” design strategy is to be able to identify and increase the priority of quasars from DESI spectra so that they will be repeatedly observed. At , the Ly forest redshifts into the DESI spectral coverage (see §II.1), making these targets especially valuable probes of large-scale structure at high redshifts. However, because we do not know in advance which quasars are at , we are forced to learn this from the DESI spectra themselves. Once Ly-forest QSOs have been identified, we update the MTL with the results to ensure that future DESI observations will observe these targets whenever possible. The MTL allows us to identify times when observations of targets failed, for example, due to being assigned to a positioner which has stopped functioning. Observations of these targets may then be repeated in the future.
The need to use information about past observations in order to inform future observations leads to the following survey operations mode or sequence:
1.
@@
The DESI spectroscopic pipeline processes each night’s observations as they come in, and usually redshifts are available the following morning (see §III.1 and Guy et al., 2023).999The DESI team informally and fondly refers to this process as “redshifts by breakfast.” Following a quality assurance process by members of the survey operations team, the reduced data are incorporated into the MTL, roughly twice per week (Schlafly et al., 2023).
Tiles are selected for observation each night automatically by the Next Tile Selector (NTS; Schlafly et al., 2023). The basic scheme is to prefer equatorial fields while minimizing slewing and obtaining the minimum airmass possible. The NTS avoids observing areas of the sky where observations have been made that have not been incorporated into the MTL, so that any remaining Ly quasars in those regions can be identified. The NTS is also responsible for deciding whether to observe bright, dark, or backup tiles. This decision is made on the basis of the current survey speed, which is based on real-time measurements of the transparency, seeing, and sky background from the Exposure Time Calculator (ETC; Kirkby et al., 2026). The survey speed is expressed as a fraction relative to nominal dark conditions, where the sky background is 21.07 mag arcsec-2 in the -band, the seeing is 11, and conditions are photometric. When the survey speed is greater than 0.4 we observe the dark program; when the survey speed is greater than 0.08 but less than 0.4 we observe the bright program; otherwise we observe backup program tiles.
@@
II.3.2 Survey Validation
Although the primary focus of DR1 is on the new main-survey data, DR1 also includes all the SV data which were taken before the start of the main survey:
1.
SV1 (Target Selection Validation) tiles were used to verify and refine the target-selection algorithms for the main survey, and include dedicated secondary-target tiles which were used for special programs before we developed the special survey (see §II.3.3).
2.
@@ II.3.3 Special Tiles
III Data Reduction and Data Products
In this section we describe the major DR1 data products and how they are produced. First, in §III.1 we present a brief overview of how the two-dimensional spectra are reduced to derive redshifts, redshift-quality flags, and spectral classifications. Next, in §III.2 we summarize the number of unique redshifts in DR1 and show the sample distributions in redshift and on the sky. In §III.3 we outline the contents and organization of the data products included in DR1, and in §III.4 we provide a brief introduction to the large-scale structure (LSS) catalogs used to carry out the DESI cosmological analysis and describe where and how those catalogs can be accessed. Finally, in §III.5 we describe the current set of value-added catalogs (VACs) accompanying DR1.
III.1 From Raw Spectra to Redshifts
The following three subsections describe the data-reduction pipeline (§III.1.1), the determination of redshifts and spectral classifications (§III.1.2), and the improvements in the spectroscopic pipeline relative to the data released in the EDR (§III.1.3).
III.1.1 Spectral Extraction and Calibration
@@
Using these data, we derive a master bias frame for each CCD, the spectral trace coordinates, wavelength calibration parameters, two-dimensional point-spread functions (PSFs), and a flat-field correction for each fiber of each camera. The processing of a scientific exposure acquired during the subsequent night proceeds as follows: We first pre-process each CCD image by subtracting the bias, the overscan level, and the dark current. We mask pixels affected by cosmic-ray hits and CCD defects, convert counts into electrons, and estimate the noise per pixel. We then adjust the coordinates of the spectral traces and the wavelength calibration of each fiber using the known wavelengths of the emission lines from the night-sky spectrum, while offsetting this solution such that the extracted spectra will be in a vacuum Solar System barycentric frame. These coordinate corrections are typically smaller than a CCD pixel relative to the afternoon calibrations. Next, we extract one-dimensional spectra using the spectroperfectionism algorithm of bolton10a. The results are counts per fiber and wavelength on a common wavelength grid (vacuum, barycentric), along with the resolution matrix, which quantifies the effective instrument resolution of each extracted spectrum as a function of wavelength. The spectra are then flat-fielded and sky-subtracted. Next, we use observations of F-type stars to determine the instrument throughput and to convert the measured counts into calibrated spectral energy density.111111All extracted (one-dimensional) DESI spectra are in units of . Finally, depending on which data products are being generated (see §III.3), multiple observations of the target or TILEID (see §II.3) are coadded using optimal (inverse variance) weights.
@@
We record the resolution matrix for each spectrum as a band-diagonal sparse matrix. A perfect resolution model should be multiplied by this matrix to achieve the equivalent observed model at the effective resolution of DESI. The underlying mathematics and motivation for the resolution matrix are described in §3 of bolton10a and in §4.5 and Appendices C and D of Guy et al. (2023). We refer interested readers to the repository of DESI tutorials (see §IV.3) for worked examples of how to use the resolution matrix in scientific analyses.
@@
For each spectrum, DESI also measures the template signal-to-noise ratio squared (TSNR2) for each of its major target classes BGS, LRG, ELG, QSO, and LYA.121212Here, LYA refers to QSOs which have been spectroscopically confirmed to be Ly forest () quasars. Also note that we do not derive a TSNR2 value for MWS targets. These quantities measure a wavelength-averaged, squared, signal-to-noise ratio, where the noise is determined from the current observations (CCD noise, sky level), the signal amplitude depends on the flux calibration, and the variation of the signal with wavelength is designed to best predict the redshift measurement precision of a given target class. For example, TSNR2_ELG gives larger weight to redder wavelengths covering the [O II] doublet at the redshifts of typical DESI ELG targets, while TSNR2_QSO gives more weight to bluer wavelengths important for measuring QSO emission lines. TSNR2_LRG is scaled to derive the effective exposure time (described in §II.3) of DESI dark-time exposures, while TSNR2_BGS is used to derive the effective exposure time for bright-time tiles, though this is also normalized to the reference dark conditions (see §4.14 of Guy et al. 2023 for more details).
III.1.2 Redshifts and Spectral Classifications
@@
With the fully calibrated spectra in-hand, we use Redrock to determine the optimum redshift and spectral classification of each object (ross20a; anand24a, Bailey et al. 2025, in preparation). Briefly, Redrock fits principal component analysis (PCA) templates of three broad, independent classes of objects—stars, galaxies, and quasars, corresponding to spectral type STAR, GALAXY, and QSO, respectively—on a spectral type-dependent grid of redshift. The fit which produces the lowest overall value yields the best-fitting redshift (Z), redshift uncertainty (ZERR), spectral type (SPECTYPE), and template coefficients (COEFF). In addition, Redrock reports the (minimum) of the best fit (CHI2) and (DELTACHI2), which is the difference between the two lowest values across all three spectral classes. Interpreted using Gaussian statistics, DELTACHI2 represents the statistical significance of the best fit relative to the next best fit; for example, a DELTACHI2 value of 25 implies that the best fit is “better” than the next best fit (not accounting for systematic uncertainties). Finally, besides redshifts and classification, a key metric reported by Redrock is the per-object redshift warning bitmask, ZWARN. In essence, a value of ZWARN=0 indicates that there are no known problems with either the input spectroscopic data nor with the corresponding redshift fit. We refer the reader to the DESI data model131313https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#zwarn, which documents the full set of possible ZWARN bits and their definitions, and to §3.2 of DESI Collaboration et al. (2024b) for more discussion.
@@
Finally, the DESI pipeline carries out three additional processing steps: emission-line fitting (emlinefit; Raichoor et al., 2023); a fit to the Mg II doublet (Chaussidon et al., 2023); and a quasar neural-network classifier called QuasarNET (busca18a; farr20a; green25a), all three of which are described in more detail in §3.1.4 of DESI Collaboration et al. (2024b).141414Colloquially, DESI team members refer to these steps of the pipeline as “after-burners,” since they are collaborator-contributed algorithms executed after the original core pipeline Redrock results have been written out. Detailed analyses show that these after-burners are essential for evaluating successful ELG redshifts and for correctly classifying and measuring the correct redshifts for some classes of QSOs (Alexander et al., 2023; Lan et al., 2023; Raichoor et al., 2023; Chaussidon et al., 2023; DESI Collaboration et al., 2024a), so the after-burner catalogs are included as standard pipeline outputs.
| Parameter | Value | ReferenceaaReferences G23, DESI24c, and K24 are Guy et al. (2023), DESI Collaboration et al. (2024d), and koposov24a, respectively. |
@@
III.1.3 Pipeline Improvements and Algorithmic Performance
Algorithmic updates for the Iron production in DR1 (relative to the Fuji specprod used in the EDR and Guadalupe; see §I.2) include new QSO templates split by redshift range (Brodzeller et al., 2023); improved sky subtraction by modeling the measured fiber throughput as a function of positioner location151515https://github.com/desihub/desispec/pull/1801 which replaces the simpler per-fiber, per-exposure measured normalization described in §4.7.2 of Guy et al. (2023); and multiple smaller bug-fixes. In addition, compared to the Fuji release in the EDR, most intermediate pipeline files are now gzipped to save disk space, although the coadded spectra and measured catalog files remain uncompressed for faster read-access (see §III.3).
@@ III.1.3 Pipeline Improvements and Algorithmic Performance
III.2 Sample Distributions
After just 13 months of science operations, DESI has measured confident redshifts for approximately 18.7M unique science targets across all surveys and programs, making DR1 the largest sample of extragalactic redshifts ever assembled. In the main survey alone, DR1 includes a total of 8.5M, 9.0M, and 1.2M objects with reliable redshifts from the bright, dark, and backup programs, respectively. Here and in the rest of this section, we use the (logical “and”) condition to select a sample of unique objects with confident redshifts. The criterion identifies objects with no known hardware, observing, data quality, or redshift-fitting flags (see §III.1), and ZCAT_PRIMARY chooses the “best” observation of objects which have been observed in different surveys or programs (see §III.3). For additional details regarding the full range of possible data-quality bit-masks, see the DESI data model.161616https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#bit-masks-in-desi Alternatively, most users of the DESI data will likely be interested in the tracer-dependent quantitative criteria used to define the samples used for DESI cosmological analyses, which are described in §4.2 of DESI Collaboration et al. (2024b) and DESI Collaboration et al. (2024d) for the EDR and DR1 cosmological results, respectively.